### v2

# Reinforcement Learning for Sequence Design Leveraging Protein Language Models

Abstract
Protein sequence design, determined by amino acid sequences, are essential to protein engineering problems in drug discovery. Prior approaches have resorted to evolutionary strategies or Monte-Carlo methods for protein design, but often fail to exploit the structure of the combinatorial search space, to generalize to unseen sequences. In the context of discrete black box optimization over large search spaces, learning a mutation policy to generate novel sequences with reinforcement learning is appealing. Recent advances in protein language models (PLMs) trained on large corpora of protein sequences offer a potential solution to this problem by scoring proteins according to their biological plausibility (such as the TM-score). In this work, we propose to use PLMs as a reward function to generate new sequences. Yet the PLM can be computationally expensive to query due to its large size. To this end, we propose an alternative paradigm where optimization can be performed on scores from a smaller proxy model that is periodically finetuned, jointly while learning the mutation policy. We perform extensive experiments on various sequence lengths to benchmark RL-based approaches, and provide comprehensive evaluations along biological plausibility and diversity of the protein. Our experimental results include favorable evaluations of the proposed sequences, along with high diversity scores, demonstrating that RL is a strong candidate for biological sequence design. Finally, we provide a modular open source implementation can be easily integrated in most RL training loops, with support for replacing the reward model with other PLMs, to spur further research in this domain. The code for all experiments is provided in the supplementary material.
1 Introduction
Proteins are indispensable for virtually every biological function, from providing structural integrity to managing immune responses and facilitating molecular transport within organisms. Despite their diverse roles, our current knowledge of proteins remains incomplete, representing merely a fraction of the vast protein landscape. Generating novel protein sequences holds promise to deepen our understanding of cellular mechanisms, physiological responses, disease pathways, and can significantly accelerate drug development (Stanton et al., 2022).
The goal of de novo protein design is to find sequences $x$ that maximize an objective function $f(x)$, scored by an oracle. This is a challenging discrete black box optimization problem over a combinatorial search space, that grows according to $\lvert\mathcal{A}\rvert^{L}$ where $L$ is the sequence length of the protein, and $\lvert\mathcal{A}\rvert=20$ is the number of most commonly appearing $\alpha$-amino acids to build proteins, although there are upto 22 amino acids in the genetic code of life (Crick, 1968; Koonin and Novozhilov, 2009). Designing proteins in this large search space has been attempted with various evolutionary methods, from directed evolution (Arnold, 1998) to adaptive greedy search like AdaLead (Sinai et al., 2020). A major caveat, however, is that these methods are computationally inefficient, find sequences in the same neighborhood, and do not scale well with increasing sequence length. In this regard, learning a sequence generation policy with reinforcement learning (RL) (Angermueller et al., 2020) offers a promising solution, with the potential to exploit the structure in the search space. Once such a policy has been trained, actions can be sampled and sequences can be generated at a fraction of the training cost and can generalize to unseen sequences, even in the absence of the reward model.
Recently, there has been massive efforts towards building large protein language models (PLMs)  (Jumper et al., 2021; Lin et al., 2023, 2022; Meier et al., 2021; Abramson et al., 2024) to address the protein folding problem, along with the prediction of scores (e.g., template modeling score) that estimate biological plausibility of a protein sequence and its corresponding folded structure. Motivated by recent research on large language models (LLMs) which provides evidence that LLMs can be repurposed as reward models (Rafailov et al., 2023), we propose to use PLMs as the reward function. In our work, we instantiate the 3B parameter version of ESMFold (Lin et al., 2023) as the oracle reward model. We choose ESMFold over other competitive PLMs such as AlphaFold2 due to its favourable inference, offering upto 60×60$\times$60 × faster inference time, which is crucial in an RL loop.
Our work takes a sequence first approach, similar to EvoDiff (Alamdari et al., 2023) and LaMBO-2 (Gruver et al., 2023). While protein structures are what ultimately determine their function, sequence determines structure. Moreover, a structure-first approach for protein design still needs to generate sequences, by solving the inverse folding problem in order to successfully synthesize the protein in the laboratory. We address the problem of protein design by learning policies that directly generate protein sequences, by learning to mutate random sequences. Our contributions are as follows:
  • The primary purpose of this paper is to investigate and benchmark the performance of RL algorithms where the reward is given by an oracle PLM, ESMFold (Lin et al., 2023) in our case.
  • Since the oracle reward model is expensive to query, we demonstrate an alternate approach that requires fewer evaluations under the oracle model, called ESM with Proxy-Finetuning (ESM-PF). ESM-PF jointly learns a proxy reward model and the sequence generation policy with results comparable to that of directly optimizing on the oracle. The proxy model is faster to query due to its smaller size, and is trained by periodically finetuning on pairs of previously generated sequences and their oracle scores $(x_{i},y_{i})$.
  • We perform a multi-objective evaluation of the proposed protein sequences in both the oracle and ESM-PF settings, across various sequence lengths, with consideration to biological plausibility and diversity of the sequences, both in the sequence and structure space.
  • We provide a modular open source implementation, that provides easy support for both the addition of other exploration algorithms (evolutionary, RL-based, generative model based), as well as replacing the ESMFold reward model with a PLM of the user’s choice (such as AlphaFold).
2 Related Work
Several methods have been explored for designing biological sequences.
Evolutionary Algorithms. Designing sequences by directed evolution (Arnold, 1998) was one of the earliest attempts to solve the protein design problem, wherein random mutations are performed on a set of sequences and the best performing ones are greedily chosen to be mutated further. AdaLead (Sinai et al., 2020) proposes selecting the sequences that are within a fraction $k$ of the best scoring sequence for further mutation. The fraction $k$ controls the greediness of the algorithm. Setting the threshold based on the score rather than the number of candidate sequences allows the algorithm to adapt to the optimization landscape; it can select many candidates in regions where the landscape is flat, thereby promoting diversity, while being able to focus on the best sequences when there are prominent peaks in the landscape. Covariance Matrix Adaptation Evolutionary Strategy (CMA-ES) (Hansen and Ostermeier, 2001) generates a population of candidate sequences from a set of parents and then selects the top $k$ sequences in a round to serve as parents the next round. The candidates are generated by sampling from a multivariate normal distribution, whose mean and covariance are obtained from the current set of parents, which gets updated at the end of every round. Proximal Exploration (PEX) (Ren et al., 2022) modifies the optimization objective to encourage selection of candidates that are close to wild type sequences. This is done by adding a regularization term that measures the edit distance of the sequence to the wild type.
Reinforcement Learning and GFlowNets. DyNAPPO (Angermueller et al., 2020), is a model-based reinforcement learning algorithm that generates protein sequences by optimizing a surrogate reward function learned through an ensemble. To promote exploration, the reward function penalizes identical sequences. On the other hand, GFlowNets were introduced as a method to sample from discrete, compositional objects by sequentially constructing the object (Bengio et al., 2021, 2023). Rather than the goal of reward maximization as in RL, GFlowNets learn to sample proportional to the reward function. This makes them well-suited for sampling from a diverse set of terminal states. For biological sequence design, GFlowNets (Jain et al., 2022, 2023) autoregressively construct the sequence by selecting one amino acid residue at a time. The methodology borrows some principles from Bayesian optimization, such as the estimation of epistemic uncertainty and the use of an acquisition function to score candidates.
Generative Models. Generative models such as discrete diffusion models, trained on large-scale protein sequence datasets, can generate novel proteins in sequence space (Alamdari et al., 2023). In contrast, diffusion (Yim et al., 2023) and flow-based models on the SE(3) manifold such as FoldFlow (Bose et al., 2024) and FoldFlow-2 (Huguet et al., 2024) can generate novel proteins in structure space but are unable to leverage the extensive sequence datasets. Methods such as RFDiffusion (Watson et al., 2023) have showed promising results employing a structure-first approach. Design by Adaptive Sampling (DbAS) (Brookes and Listgarten, 2018) trains a variational autoencoder (Kingma and Welling, 2013) in an unsupervised manner, while accounting for uncertainty in the oracle model. The follow-up work on Conditioning by Adaptive Sampling (CbAS) (Brookes et al., 2019) addresses the pathological behaviour of the oracle – wherein the oracle provides wildly unreliable scores for sequences in a neighborhood far away from where it was trained – by encoding knowledge about regions where the oracle is expected to be accurate. There are also other approaches to sequence design based on GANs (Goodfellow et al., 2014), such as activation maximization and FB-GAN for generating DNA sequences (Killoran et al., 2017; Gupta and Zou, 2018).
MCMC. Biological sequences can also be generated by sampling from energy-based models. MCMC and simulated annealing can generate sequences by optimizing energy functions that evaluate compatibility with constraints on the protein structures (Hie et al., 2022), leveraging protein language models for structure prediction. A smoothed energy function can also be learned (Frey et al., 2023), and sequences can be sampled from the smoothed data manifold using MCMC, following one-step denoising.
Bayesian Optimization (BO) is a standard approach for optimizing black box functions using a surrogate model of the objective (Močkus, 1975; Snoek et al., 2012) and typically involves maximizing an acquisition function. Many works have adapted Bayesian Optimization for sequence design (Stanton et al., 2022; Gruver et al., 2023; Swersky et al., 2020; Belanger et al., 2019). LaMBO (Stanton et al., 2022) uses a MLM style decoding guided with acquisition values to address the multi objective nature of real-world protein design. A follow-up work, LaMBO-2 (Gruver et al., 2023), argues to use diffusion optimized sampling models in a similar framework for better design.
3 Preliminaries
Reinforcement Learning
In RL, we model a decision making task through an environment that describes a Markov Decision Process (MDP) $<\mathcal{S,A,R,P}>$, where $\mathcal{S},\mathcal{A},\mathcal{R}$ and $\mathcal{P}$ represent the state space, action space, rewards and transition dynamics respectively. At every timestep, the agent receives a state $s_{t}\in\mathcal{S}$ and predicts an action $a_{t}\in\mathcal{A}$, to transition to a new state according to the transition probability $p(s_{t+1}\mid s_{t},a_{t})$, and receives a reward $r_{t}\in\mathcal{R}$. The agent chooses an action using a policy $\pi_{\theta}(a_{t}|s_{t})$, parameterized by $\theta$. The objective is to learn an optimal policy $\pi_{\theta}^{*}(a_{t}|s_{t})$ that maximises the expected sum of rewards from the environment. Concretely, optimization problem can be summarized as,
$\theta^{*}=\operatorname*{arg\,max}_{\theta}\mathbb{E}_{a_{0:T}\sim\pi_{\theta}}\left[\sum^{T}_{t=0}\gamma^{t}r_{t}\right]$
Problem Setting
We frame the problem of protein sequence design in an RL loop, where at each timestep the agent receives a sequence and predicts the mutation to be performed. The sequences are represented through one hot encodings of each character in the sequence. That is, given a sequence of length $L_{s}$ and an alphabet of size $L_{a}$, the state vector is denoted by $s\in\mathbb{R}^{L_{s}\times L_{a}}$. The action space is modelled as a discrete choice over $L_{s}\times L_{a}$ options, indicating the position of the mutation and the mutated character. The mutated sequence is then returned as the next state. The reward is the pTM score of the mutated sequence, which is obtained from either the oracle model or the proxy model. Our environment thus is completely deterministic, the state transitions are simple sequence mutations and the rewards are deterministically obtained from the models. Finally, we design the environment to be an infinite horizon task. RL training is simultaneously conducted on a batch of $100$ sequences. ESMFold is queried with the batch of all $100$ sequences and inference is run across 4 GPUs to obtain the rewards (i.e, pTM score).
Figure 1: Training policies with the a) oracle, b) proxy reward models.
4 Reinforcement Learning for Generating Protein Sequences
4.1 Approach
In this paper, we provide a modular approach to leverage existing protein language models for sequence design, in an RL framework. Our work relies on using pretrained LMs and then being able to train a RL based agent for mutating the sequences over long horizon lengths, for generating amino acid sequences. To this end, we first provide an overview of the different RL algorithms being tests for the protein design tasks. We note that even though prior works have explored this, we provide the first approach for RL based algorithms to leverage existing protein language models, such as the well known ESMFold model Lin et al. (2023), for these tasks.
Baselines: We provide an overview of the baselines that were tested on this benchmark. We experiment with both value based and policy based deep RL algorithms for learning mutation policies for protein sequences: (1) Deep Q Network (DQN) (Mnih et al., 2013) (2) Rainbow (Hessel et al., 2018) (3) Proximal Policy Optimization (PPO) (Schulman et al., 2017) (4) Soft Actor-Critic (SAC) (Haarnoja et al., 2018) for the base RL algorithms. We also try an additional exploration based baseline, namley (5) PPO with Random Network Distillation (RND) (Burda et al., 2019). We compare our approach with several works that have previously been proposed for similar tasks, namely (6) GFlowNets (Bengio et al., 2021, 2023) and (7) Metropolis Hastings MCMC with simulated annealing  (Metropolis et al., 1953; Hastings, 1970). Rather than autoregressively constructing the sequences as in Jain et al. (2022), we use a single-step GFlowNet to mutate sequences.
In our setup, starting sequences are randomly initialized and the policy is mutative. At each step, the action sampled from the policy, chooses one of $L_{s}\times L_{a}$ actions, describing the index to mutate and also the residue from the alphabet that will substitute the original amino acid. The policy is trained in one of two ways: (1) using the oracle reward model, or (2) using the proxy reward model, which is finetuned every $K=2000$ steps. Figure 1(a) and 1(b) illustrate the RL loop with the oracle reward model and proxy reward model with finetuning, respectively. For all methods, we run $100$ environments in parallel.
4.2 Learning the Reward Model
In this section, we outline our approach for learning a proxy to the reward function. As outlined in section 4.1, the RL based mutation process tries to maximize the reward function at each step. However, it is computationally expensive to evaluate the reward, especially for protein sequences of longer lengths.
Figure 2: Pearson correlation score between the proxy and oracle model. For readability, the curves are smoothed with exponential moving average.
Inspired by knowledge distillation (Hinton et al., 2015; Polino et al., 2018), we distill the 3B ESM model (teacher) into a smaller transformer model (Vaswani et al., 2017), with ${}15\times 10^{3}$ parameters, to serve as a proxy reward model. For learning the proxy reward model, we learn the reward function using sequences from the Atlas dataset (Vander Meersche et al., 2024) since it contains diverse samples. The reward model is then trained using a mean squared regression objective, by predicting the pTM scores for each sequence.
To ensure we learn a good reward model, we monitor the correlation scores between the proxy predicted scores and the oracle evaluations for each sequence. Additionally, once we have the pretrained proxy model, we can finetune the proxy within the training loop of the algorithms. Every $2000$ environment steps, we compute the proxy pTM scores for the proposed sequences, filter the top $K=100$ sequences, and finetune the proxy on the true pTM scores of these sequences from ESMFold for $50$ epochs. We demonstrate the effectiveness of the proxy finetuning procedure in Figure 2; which shows an increasing trend in Pearson’s correlation coefficient between the proxy and oracle model for sequences of length $50$.
4.3 Evaluating Protein Sequences
Evaluating the designed protein sequences is inherently multi-objective. In our experiments, evaluations are primarily along two axes, scoring the sequences according to biological plausibility and diversity. Biological plausibility is the useful to assess the viability of the proposed sequences within biological contexts. To this end, we report the mean predicted Template Modeling (pTM) score and the mean predicted Local Distance Difference Test (pLDDT) score across the proposed sequences:
TM score (Zhang and Skolnick, 2004).
A metric in $(0,1]$ that evaluates the structural similarity between two proteins, with higher TM-scores indicating better structural similarity. It is a variation of the Levitt-Gerstein (LG) score, which weights shorter distances between residues more strongly than longer distances. The TM-score between a target structure $s_{\text{tgt}}$ and template structure $s_{\text{tmp}}$ is defined as:
$d_{\text{TM}}(s_{\text{tgt}},s_{\text{tmp}})=\max\bigg{[}\frac{1}{L_{\text{tgt}}}\sum\limits_{i=1}^{L}\frac{1}{1+\big{(}\frac{d_{i}}{d_{0}(L_{\text{tgt}})}\big{)}^{2}}\bigg{]}$ (1)
where $L_{\text{tgt}}$ is amino acid sequence length of the target protein, $L$ is the number of residues that appear in both template and target structures, $d_{i}$ is the distance between the $i^{\text{th}}$ residues in template and target structures, and $d_{0}(L_{\text{tgt}})$ is taken as $1.24\sqrt[3]{(L_{\text{tgt}}-15})-1.8$ and normalizes sequence length
lDDT score (Mariani et al., 2013).
A metric in $(0,100]$ which provides a per-residue measure of confidence in the local accuracy of a predicted protein structure. lDDT estimates how well the prediction would agree with an experimental structure and a score of above 70 usually corresponds to a correct backbone prediction with misplacement of some side chains. An lDDT score of 90 or greater indicates high confidence in the predicted structure, while a score below 50 indicates poor confidence.
In addition to obtaining high-scoring sequences, we are interested in obtaining diverse sequences. Evaluating diversity is multi-dimensional and can be viewed through different lenses such as diversity in sequence space, structural diversity of the protein, number of modes discovered, and hyper-volume. In our experiments, we report multiple metrics for a comprehensive evaluation of sequence diversity through all these lenses:
Diversity in Sequence Space.
The Mean Pairwise Hamming Distance (MP-HD) over a dataset $\mathcal{D}$ of sequences is a standard metric to evaluate the diversity in the sequence space (Angermueller et al., 2020; Jain et al., 2022), and is measured as:
[…]
where $d_{\text{HD}}(.,.)$ is the hamming distance between two sequences.
Structural Diversity.
We first obtain the structures $\mathcal{S}$ from $\mathcal{D}$ by performing a forward pass through a protein structure prediction model (ESMFold in our case) and then compute the Mean Pairwise TM score (MP-TM) and Mean Pairwise RMSD (MP-RMSD).
[…]
Since TM-score is a closeness metric and RMSD is a distance metric, we typically want low MP-TM ($\downarrow$) and high MP-RMSD ($\uparrow$).
However, the reward models could be imperfect or misspecified. To quantify this, we compare outputs of ESMFold with that of AlphaFold. Concretely, we compare the mean pTM and pLDDT scores predicted by each model across the proposed sequences for each algorithm. We also report the mean RMSD between the structure predicted by both models, over all proposed sequences.
5 Experiments and Analysis
In this section, we outline our experiments for sequence design. In all our experiments, we use ESMFold either directly as the reward model or as a signal to finetune the proxy model. Our main goal is to test the efficacy of different sequence design algorithms. Through our experiments, we analyze the performance of all methods and ask the following questions.
Figure 3: Learning curves for optimizing the reward (pTM) from ESMFold on sequences of length 50. The x-axis is the number of queries to the reward model.
How effectively do various sequence design algorithms optimize the reward from ESMFold?
Table 1 and 4 (appendix) present the results for all methods evaluated in the infinite horizon setting for sequences of lengths $50$ and $100$, respectively. The reward used in these experiments is the pTM scores obtained from the oracle model (ESMFold). All results are averaged over three random seeds, and all experiments for this analysis were run on 4 A100 GPUs. The results in these tables highlight the effectiveness of RL algorithms and GFlowNets for sequence design. Figure 3 illustrates the training curves for optimizing the oracle reward for sequences of length $50$.
Our training results show MCMC outperforms PPO in terms of the pTM score. While this result might suggest that MCMC is superior for sequence design, it is important to consider the broader context and practical implications.
Table 1: Optimizing on pTM scores from ESMFold (Sequence length $50$): Final scores of the generated protein sequences with a budget of $5\times 10^{7}$ oracle reward evaluations. Best protein confidence scores and second best diversity scores are highlighted. All scores are computed on the outputs of the 3B ESMFold model.
Protein Confidence Diversity Score pTM ($\uparrow$) pLDDT ($\uparrow$) MP-TM ($\downarrow$) MP-RMSD ($\uparrow$) MP-HD ($\uparrow$) Random 0.43$\pm$0.04 63.0$\pm$6.21 0.52$\pm$0.08 2.04$\pm$0.35 0.88$\pm$0.01 MH-MCMC 0.86$\pm$0.01 92.9$\pm$0.17 0.71$\pm$0.15 1.17$\pm$0.52 0.74$\pm$0.04 DQN 0.35$\pm$0.04 56.14$\pm$3.53 0.83$\pm$0.15 0.93$\pm$0.69 0.03$\pm$0.03 SAC 0.65$\pm$0.01 85.1$\pm$0.47 0.88$\pm$0.05 0.72$\pm$0.08 0.79$\pm$0.01 PPO 0.81$\pm$0.02 91.4$\pm$2.05 0.99$\pm$0.01 0.21$\pm$0.33 0.03$\pm$0.05 PPO-RND 0.64$\pm$0.01 84.43$\pm$0.75 0.91$\pm$0.05 0.70$\pm$0.09 0.60$\pm$0.04 Rainbow 0.55$\pm$0.03 77.7$\pm$3.56 0.79$\pm$0.06 1.22$\pm$0.17 0.70$\pm$0.02 GFN $(\beta=6)$ 0.52$\pm$0.01 74.0$\pm$1.00 0.69$\pm$0.03 1.39$\pm$0.05 0.87$\pm$0.01 GFN $(\beta=32)$ 0.62$\pm$0.02 82.5$\pm$0.73 0.88$\pm$0.02 0.74$\pm$0.09 0.84$\pm$0.02 GFN $(\beta=64)$ 0.63$\pm$0.01 83.7$\pm$0.55 0.90$\pm$0.02 0.76$\pm$0.04 0.82$\pm$0.01
In contrast, learning a policy, such as with RL or GFlowNets, offers several compelling advantages:
  1. 1.
    Generalization: Once a policy is trained, it can generalize to new unseen sequences without requiring retraining from scratch. This ability to adapt to new sequences without starting over is crucial for efficiency.
  2. 2.
    Cost efficient at evaluation time: After training, the policy can generate multiple candidates simply by sampling actions from the policy. This process does not require access to the reward model and incurs only a fraction of the original training cost, making it much more efficient for large-scale sequence generation.
  3. 3.
    Adaptability: The policy can be finetuned with additional data, allowing for iterative improvements and adaptations to new requirements or environments. This flexibility is not easily achievable with MCMC.
Although MCMC shows better performance in terms of the pTM score, it has significant limitations. Specifically, MCMC lacks the ability to generalize to unseen sequences and must be run from scratch for each new sequence. This leads to increased computational costs and time for every new instance.
Can a smaller, cost-efficient proxy model be trained to approximate ESMFold scores accurately? Can sequences be generated by optimizing the rewards from this proxy model?
We find that all RL methods and GFlowNets are able to maintain high oracle pTM scores by optimizing on the proxy pTM (Fig  4 and Table 3), while cutting down the runtime by an order of magnitude (Table  5). All results are averaged over three random seeds, and all experiments for this analysis were run on a single RTX8000 GPU. We present the complete results in Table 3.
Most notably, while MCMC demonstrated the best performance when optimizing directly on the oracle (Tab. 1), it tends to overcommit to the proxy scores, ultimately compromising its performance on the oracle pTM (Fig  4). We hypothesize that MCMC, being a hill-climbing method, may become trapped in local optima that are favorable according to the proxy model but not according to the oracle. Given that the proxy is a smaller, less complex model, its reward landscape likely contains different local optima that do not align with those of the oracle.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 4: Learning curves for optimizing the reward from the proxy model on sequence length $50$. Both the proxy pTM and oracle pTM are shown, with the x-axis referring to number of proxy reward queries.
Table 2: Comparing A100 GPU runtimes, in minutes, for oracle and proxy-finetune experiments (Sequence length $50$). Proxy runtimes are faster by an order of magnitude.
Oracle Runtime Proxy Runtime PPO $1200$ $75$ PPO-RND $1270$ $80$ DQN $2400$ $345$ SAC $2640$ $285$ Rainbow $2760$ $300$ GFN $988$ $75$
Table 3: Optimizing on pTM scores from the finetuned proxy model (ESM-PF; Sequence length $50$): Final scores of the generated protein sequences with a budget of $5\times 10^{7}$ proxy reward evaluations. Best protein confidence scores and second best diversity scores are highlighted. All scores are computed on the outputs of the 3B ESMFold model.
Protein Confidence Diversity Score pTM ($\uparrow$) pLDDT ($\uparrow$) MP-TM ($\downarrow$) MP-RMSD ($\uparrow$) MP-HD ($\uparrow$) Random 0.32$\pm$0.03 54.0$\pm$2.9 0.45$\pm$0.02 2.31$\pm$0.12 0.84$\pm$0.00 MH-MCMC 0.25$\pm$0.03 48.6$\pm$1.8 0.52$\pm$0.06 2.11$\pm$0.34 0.76$\pm$0.01 DQN 0.32$\pm$0.14 54.8$\pm$10.3 0.81$\pm$0.05 1.06$\pm$0.11 0.04$\pm$0.02 SAC 0.65$\pm$0.04 84.7$\pm$1.5 0.93$\pm$0.02 0.54$\pm$0.06 0.67$\pm$0.09 PPO 0.80$\pm$0.04 88.3$\pm$0.7 0.99$\pm$0.00 0.12$\pm$0.07 0.05$\pm$0.06 PPO-RND 0.58$\pm$0.04 80.7$\pm$1.8 0.85$\pm$0.08 0.83$\pm$0.13 0.60$\pm$0.01 Rainbow 0.51$\pm$0.08 72.3$\pm$7.6 0.69$\pm$0.09 1.39$\pm$0.50 0.48$\pm$0.02 GFN $(\beta=6)$ 0.36$\pm$0.01 57.1$\pm$1.0 0.51$\pm$0.06 1.94$\pm$0.30 0.85$\pm$0.02 GFN $(\beta=32)$ 0.56$\pm$0.01 77.1$\pm$1.1 0.80$\pm$0.04 0.99$\pm$0.24 0.82$\pm$0.00 GFN $(\beta=64)$ 0.61$\pm$0.00 81.9$\pm$0.6 0.93$\pm$0.02 0.58$\pm$0.10 0.79$\pm$0.02
How do different algorithms compare when holistically evaluated on biological plausibility and diversity scores?
In both experiments, whether optimizing on the oracle or the proxy reward model, the generated sequences are evaluated using five metrics: pTM and pLDDT scores for biological plausibility, two structural diversity metrics (MP-TM and MP-RMSD), and MP-HD, which measures diversity in the sequence space. Table 1 and 3 summarizes all five metrics for both settings.
Since the evaluation of the generated sequences is inherently multi-objective, we visualize Pareto plots, combining one biological plausibility metric (in this case, the pTM score) with each of the three diversity metrics. Figures 5 and 7 provide these visualizations, highlighting only those methods that achieve a pTM score of at least $0.5$.
Figure 5: Pareto plots for optimization on the finetuned proxy scores: Plotted across two structural diversity metrics and one sequence diversity metric (MP-HD). Only methods with atleast $0.5$ pTM are highlighted. PPO, SAC, and GFlowNets form the Pareto front across all 3 diversity metrics.
From Figure 5 it is evident that PPO, SAC, and GFN ($\beta=32,64$) are the top-performing methods, establishing the Pareto front across all three diversity metrics. Notably, Rainbow also appears on the Pareto front for structural diversity metrics such as MP-TM and MP-RMSD, although it does not achieve this for MP-HD. Additionally, GFlowNets exhibit flexibility in balancing pTM scores with diversity. By adjusting the hyperparameter $\beta$, GFlowNets can attain different positions on the Pareto front, as highlighted by the points marked with an "x" in Fig 5).
While MCMC remains in the Pareto set for sequences of length 50, demonstrating good diversity and bio-plausibility scores, we observe a decline in its diversity as we scale to longer sequences. For sequences of length 100, PPO becomes comparable to MCMC in terms of both bio-plausibility scores and diversity (Table 4 on sequence length $100$).
How does the setting (infinite horizon vs finite horizon) affect performance?
Figure 6: Effect of episode length in the finite horizon setting on the pTM-diversity tradeoff. The ablation is performed for PPO trained on the ESM-PF proxy model for sequences of length $50$.
Here, we perform an ablation study on episode length in the finite horizon setting, contrasting it with our previous results in the infinite horizon setting. Due to brevity and computational limitations, we conduct experiments solely on PPO optimized with ESM-PF on sequences of length $50$. We examine different episode lengths (i.e., $L_{s}$ and $5L_{s}$) and observe their effects on the oracle pTM score and diversity metric. For conciseness, we consider only the MP-HD for diversity. Generally, for the same reward budget, shorter episode lengths result in low pTM and high diversity, whereas longer episode lengths result in high pTM and low diversity (Figure 6). Depending on the confidence in the reward model, the RL loop can be run in different settings to trade off diversity (accounting for reward misspecification) for bio-plausibility scores, and vice-versa.
6 Limitations
Our results are constrained by sequence length due to the prohibitive computational requirements associated with longer sequences. The findings presented in this paper rely on either the proxy or the 3B ESMFold oracle model. However, the reward model may itself be uncertain or misspecified. Future research should consider extending these results to optimization on other PLMs, such as AlphaFold2, or the larger 15B parameter ESMFold model. Additionally, exploring larger proxy reward models could provide a more accurate approximation of the oracle reward model for longer sequences. Though we do not foresee any negative consequences arising from this work, PLMs could be used maliciously by users.
7 Conclusion
In this work, we propose to leverage existing pretrained PLMs to learn a mutation policy for generating protein sequences and demonstrate the efficacy of this approach. We benchmark the performance of the algorithms in a wide range of settings – training with the $3$B ESMFold and the proxy, generating various sequence lengths, and studying the pTM-diversity tradeoff for different episode lengths. To this end, we demonstrate and compare performance, of several existing deep RL algorithms compared to well known baselines, and show that RL can be quite effective in this domain. We share a modular implementation in the appendix, with support for adding a PLM of the user’s choice as the reward model, as well as adding other RL algorithms.
Acknowledgements
The authors would like to thank Compute Canada and Mila for the computing resources required for this work. JS would like to thank Kausthubh Ramachandran, Michael Jendrusch, Moksh Jain and Nanda Harishankar Krishna for fruitful discussions and for providing feedback on initial drafts of the paper. DN and SEK would also like to acknowledge CIFAR.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We show how PLMs can be leveraged for protein sequence design in an RL paradigm and provide comprehensive evaluation of RL baselines across both sequence-based and structural metrics.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a section on the limitations of the current work and potential extentions to address these limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the code for the RL baselines and environment implementations used in the paper. The repositories include instructions on how to reproduce all the results in the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code to reproduce the experiments has been added to the supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental section contains the problem setting and a description of how the code can be used.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We run all experiments for multiple seeds and report the mean and standard deviation across runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include the number of GPUs used per experiment and the average running time.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, all authors have reviewed the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We address potential harms in the limitations section.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The work in this paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the papers that introduce the PLMs used in this paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We document the code released in a README.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
Appendix A Additional Figures
Figure 7: Pareto plots for optimizing proteins of length $50$ on ESMFold pTM scores: Plotted across two structural diversity metrics and one sequence diversity metric (MP-HD). Only methods with atleast $0.5$ pTM are highlighted.
Appendix B Experiments on Longer Sequence Lengths
Table 4: Optimizing on pTM scores from ESMFold (Sequence length $100$): Final scores of the generated protein sequences with a budget of $10\times 10^{7}$ oracle reward evaluations. Best protein confidence scores and second best diversity scores are highlighted. All scores are computed on the outputs of the 3B ESMFold model.
Protein Confidence Diversity Score pTM ($\uparrow$) pLDDT ($\uparrow$) MP-TM ($\downarrow$) MP-RMSD ($\uparrow$) MP-HD ($\uparrow$) Random 0.40$\pm$0.00 46.9$\pm$0.48 0.35$\pm$0.01 3.99$\pm$0.02 0.90$\pm$0.02 MH-MCMC 0.93$\pm$0.00 92.2$\pm$1.09 1.00$\pm$0.00 0.04$\pm$0.03 0.01$\pm$0.01 DQN 0.37$\pm$0.03 42.2$\pm$2.74 0.62$\pm$0.14 2.76$\pm$0.75 0.15$\pm$0.12 SAC 0.71$\pm$0.02 78.5$\pm$2.32 0.71$\pm$0.12 1.87$\pm$0.43 0.66$\pm$0.06 PPO 0.86$\pm$0.02 88.9$\pm$1.90 1.00$\pm$0.00 0.03$\pm$0.03 0.01$\pm$0.01 PPO-RND 0.50$\pm$0.05 55.4$\pm$5.16 0.91$\pm$0.11 0.53$\pm$0.64 0.16$\pm$0.20 Rainbow 0.45$\pm$0.06 52.7$\pm$3.70 1.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 GFN $(\beta=6)$ 0.42$\pm$0.03 48.6$\pm$1.13 0.35$\pm$0.01 3.99$\pm$0.11 0.90$\pm$0.01 GFN $(\beta=32)$ 0.46$\pm$0.01 53.7$\pm$1.39 0.38$\pm$0.01 3.59$\pm$0.10 0.88$\pm$0.02 GFN $(\beta=64)$ 0.55$\pm$0.04 63.6$\pm$1.87 0.55$\pm$0.03 2.66$\pm$0.33 0.78$\pm$0.07
Table 5: Optimizing on pTM scores from ESMFold (Sequence length $150$): Final scores of the generated protein sequences with a budget of $10\times 10^{7}$ oracle reward evaluations. Best protein confidence scores and diversity scores are highlighted. All scores are computed on the outputs of the 3B ESMFold model.
Protein Confidence Diversity Score pTM ($\uparrow$) pLDDT ($\uparrow$) MP-TM ($\downarrow$) MP-RMSD ($\uparrow$) MP-HD ($\uparrow$) MH-MCMC 0.93$\pm$0.00 89.77$\pm$0.60 1.00$\pm$0.00 0.03$\pm$0.02 0.01$\pm$0.00 SAC 0.36$\pm$0.02 38.22$\pm$0.93 0.34$\pm$0.01 4.81$\pm$0.12 0.87$\pm$0.02 PPO 0.91$\pm$0.01 87.64$\pm$1.55 1.00$\pm$0.00 0.11$\pm$0.04 0.03$\pm$0.01 PPO-RND 0.39$\pm$0.02 43.65$\pm$5.36 0.83$\pm$0.01 1.24$\pm$0.18 0.24$\pm$0.03 GFN $(\beta=6)$ 0.35$\pm$0.01 37.70$\pm$1.96 0.34$\pm$0.02 4.74$\pm$0.21 0.86$\pm$0.03 GFN $(\beta=32)$ 0.36$\pm$0.01 39.97$\pm$2.23 0.35$\pm$0.03 4.67$\pm$0.25 0.87$\pm$0.02 GFN $(\beta=64)$ 0.40$\pm$0.03 42.04$\pm$1.07 0.62$\pm$0.01 2.62$\pm$2.15 0.52$\pm$0.46
Appendix C Biophysical Measures
In this section, we present several biophysical properties computed from the amino acid sequences of the generated proteins using Biopython [Cock et al., 2009]. Namely, we report the molecular weight ($W_{mol}$), instability index, isoelectric point, grand average of hydropathy (GRAVY) computed with the Kyte-Doolittle scale Kyte and Doolittle [1982]. These properties are then compared to a reference distribution defined by naturally occurring proteins from the Protein Data Bank (PDB) [Burley et al., 2018] using the Distributional Conformity Score (DCS) introduced in Frey et al. [2023]. Tables 6, 7, and 8 report the metrics for proteins of residue lengths $50,100$ and $150$ respectively.
Table 6: Biophysical metrics of $50$-residue proteins generated by optimizing pTM scores using ESMFold. The last column is the Distributional Conformity Score comparing generated proteins to reference proteins of the same length in PDB.
$W_{mol}$ Instability Index Isoelectric Point (pI) GRAVY DCS ($\uparrow$) MH-MCMC $5598.95$ $35.90$ $5.87$ $0.22$ $0.53$ DQN $5990.59$ $41.57$ $7.69$ $-0.16$ $0.31$ SAC $6066.13$ $47.09$ $7.75$ $0.30$ $0.39$ PPO $5955.30$ $49.38$ $6.33$ $0.57$ $0.56$ PPO-RND $6078.18$ $54.13$ $8.11$ $0.18$ $0.41$ Rainbow $5908.24$ $41.17$ $6.61$ $0.37$ $0.47$ GFN $(\beta=6)$ $6166.45$ $44.34$ $8.04$ $-0.36$ $0.29$ GFN $(\beta=32)$ $6106.51$ $51.47$ $9.18$ $-0.08$ $0.41$ GFN $(\beta=64)$ $6071.02$ $54.17$ $8.25$ $-0.29$ $0.41$
Table 7: Biophysical metrics of $100$-residue proteins generated by optimizing pTM scores using ESMFold. The last column is the Distributional Conformity Score comparing generated proteins to reference proteins of the same length in PDB.
$W_{mol}$ Instability Index Isoelectric Point (pI) GRAVY DCS ($\uparrow$) MH-MCMC $11470.28$ $35.67$ $5.52$ $-0.21$ $0.53$ DQN $12096.87$ $43.84$ $8.74$ $-0.54$ $0.29$ SAC $12079.58$ $46.09$ $6.78$ $0.87$ $0.27$ PPO $11697.12$ $34.69$ $10.55$ $0.36$ $0.52$ PPO-RND $12009.17$ $51.34$ $7.10$ $-0.52$ $0.33$ Rainbow $12121.63$ $51.81$ $9.25$ $-0.33$ $0.17$ GFN $(\beta=6)$ $12058.39$ $44.49$ $7.55$ $-0.46$ $0.34$ GFN $(\beta=32)$ $12091.40$ $44.00$ $7.47$ $-0.35$ $0.29$ GFN $(\beta=64)$ $12228.52$ $45.64$ $8.23$ $-0.14$ $0.18$
Table 8: Biophysical metrics of $150$-residue proteins generated by optimizing pTM scores using ESMFold. The last column is the Distributional Conformity Score comparing generated proteins to reference proteins of the same length in PDB.
$W_{mol}$ Instability Index Isoelectric Point (pI) GRAVY DCS ($\uparrow$) MH-MCMC $17818.09$ $32.19$ $7.99$ $-0.16$ $0.09$ SAC $17841.44$ $41.99$ $7.17$ $-0.47$ $0.22$ PPO $17971.68$ $48.31$ $7.53$ $0.09$ $0.03$ PPO-RND $17920.13$ $50.46$ $8.87$ $-0.37$ $0.19$ GFN $(\beta=6)$ $18031.62$ $41.95$ $7.15$ $-0.45$ $0.16$ GFN $(\beta=32)$ $18008.17$ $46.20$ $7.49$ $-0.39$ $0.15$ GFN $(\beta=64)$ $18129.90$ $38.15$ $7.19$ $-0.53$ $0.09$
Table 9: Optimizing on pTM scores from finetuned proxy model (ESM-PF): Distributional Conformity Score
DCS $\uparrow$
$L=50$ $L=100$
PPO 0.3074 0.1163
SAC 0.3361 0.3280
DQN 0.4807 0.1397
PPO-RND 0.3268 0.0457
Rainbow 0.4340 0.6140
GFN $(\beta=6)$ 0.3622 0.2395
GFN $(\beta=32)$ 0.3667 0.3803
GFN $(\beta=64)$ 0.2908 0.2376
MH-MCMC 0.2583 0.0321
Table 10: Optimizing on pTM scores from ESMFold: Mean absolute error in normalized frequency of amino acids between proposed sequences and naturally occurring proteins from the PDB.
MAE $\downarrow$
$L=50$ $L=100$ $L=150$
PPO 0.047 0.029 0.025
SAC 0.036 0.036 0.021
DQN 0.027 0.022 0.019
PPO-RND 0.032 0.017 0.021
Rainbow 0.024 0.018 0.050
GFN $(\beta=6)$ 0.021 0.016 0.023
GFN $(\beta=32)$ 0.025 0.018 0.022
GFN $(\beta=64)$ 0.026 0.021 0.021
MH-MCMC 0.031 0.017 0.023
Table 11: Optimizing on pTM scores from finetuned proxy model (ESM-PF): Mean absolute error in normalized frequency of amino acids between proposed sequences and naturally occurring proteins from PDB
MAE $\downarrow$
$L=50$ $L=100$
PPO 0.0526 0.0285
SAC 0.0462 0.0390
DQN 0.0207 0.0230
PPO-RND 0.0307 0.0204
Rainbow 0.0321 0.0265
GFN $(\beta=6)$ 0.0228 0.0195
GFN $(\beta=32)$ 0.0275 0.0233
GFN $(\beta=64)$ 0.0363 0.0212
MH-MCMC 0.0203 0.0240